OpenStudy (anonymous):

(4,64*х)/(4,64+х)=4,41

OpenStudy (anonymous):

@Jamierox4ev3r

HanAkoSolo (jamierox4ev3r):

why are there commas in this?

HanAkoSolo (jamierox4ev3r):

is this what you mean : \(\frac{4.64*x}{4.64*x}\)=4.41

OpenStudy (anonymous):

x=88.967 i found the answer i need explanation

OpenStudy (anonymous):

@Jamierox4ev3r

hero (hero):

\[\frac{4.64x}{4.64 + x} = 4.41\] Multiply both sides by 4.64 + x to get 4.64x = 4.41(4.64 + x) Distribute 4.41 over 4.64 + x: 4.64x = 4.41*4.64 + 4.41x 4.64x = 20.4624 + 4.41x Subtract 4.41x from both sides to place like terms on the same side: 4.64x - 4.41x = 20.4624 x is common to both terms on the left side, so factor it out: (4.64 - 4.41)x = 20.4624 Subtract: .23x = 20.4624 Divide both sides by .23 to isolate x: x = 20.4624/.23 x = 88.967

hero (hero):

@Mr.ClayLordMath, Did that help?

OpenStudy (anonymous):

ofc thnx bro your best math solver :)